Friendly’s Quick Loans, Inc., offers you “six for seven or I knock on your door.” This means you get $6 today and repay $7 when you get your paycheck in one week (or else).

a. If you were brave enough to ask, what APR would Friendly’s say you were paying? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What’s the EAR Friendly’s earns on this lending business? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

7=6*(1+r/100)

(7/6)=(1+r/100)

(1+r/100)=1.1667

r=1.1667-1

=16.67%(Approx).

a.APR=16.67%*52 weeks

=866.67%(Approx).

b.EAR=[(1+APR/m)^m]-1
where m=compounding periods

=[(1+0.1667)^52]-1

=302,753.82%(Approx).